# An Efficient and Adaptive Next Edit Suggestion Framework with Zero Human Instructions in IDEs

### v1

@@ 4.2. Results
Table 3. Comparison of location prediction performance with varying maximum edit history lengths on the Qwen3-4B+SFT model
@@ 4.2.1. RQ1 NES-Location:
Impact of SFT: Before post-training, the best base model, Claude 4, achieves 67.0% and 9.6% average accuracy on the -do and -keep tasks, respectively. Models with small size, such as Qwen2.5-Coder-7B, perform poorly, with only 10.1% and 3.7% on the -do and -keep tasks, respectively. NES SFT dataset significantly improves models with small size. For -do, Qwen2.5-Coder-7B jumps from 10.1% to 62.8%, Seed-Coder-8B-Instruc from 9.3% to 72.6%, and Qwen3-4B from 6.6% to 72.6%. For -keep, Qwen2.5-Coder-7B rises from 3.7% to 94.0%, Seed-Coder-8B-Instruc from 2.9% to 94.1%, and Qwen3-4B from 1.9% to 89.4%.
Due to the limits of low-latency requirements and training resource constraints, large models were not trained on the SFT dataset. Additionally, Zeta does not provide navigation functionality, so its results are excluded from Table 2.
RQ1 Answer about SFT: The NES SFT dataset significantly boosts model performance in predicting the next editing position, improving Qwen3-4B’s accuracy from 6.6% to 72.6% on the -do task and from 1.9% to 89.4% on the -keep task.
Impact of DAPO: Models with smaller sizes trained on the DAPO dataset achieve slight improvements in average navigation prediction (-do) accuracy: Qwen2.5-Coder-7B improves from 62.8% to 65.6%, Seed-Coder-8B-Instruc increases from 72.6% to 77.7%, and Qwen3-4B rises from 72.6% to 75.6%. However, on the -keep task, Qwen2.5-Coder-7B declines from 94.0% to 86.5%, Seed-Coder-8B-Instruct declines from 94.1% to 85.0%, and Qwen3-4B drops from 89.4% to 81.6%.
The DAPO-trained model tends to suggest location changes more frequently. Because changing is more frequently in the real-world development practices, We accept the minor drop in -keep performance in favor of improved -do accuracy. The NES-Location Model aims to streamline navigation, so we’ve adopted DAPO, prioritizing the higher success rate in predicting necessary jumps over avoiding suggestions, as it better supports developers’ workflow.
RQ1 Answer about DAPO: The DAPO-trained model tends to suggest location changes more frequently, leading to a slight drop in -keep task accuracy. However, we have decided to adopt the DAPO-trained NES-Location model, which demonstrates better performance on the -do task, as location changes are more prevalent in real-world development practices.
Impact of Edit History Length: Table 3 demonstrates the impact of historical context length (Max Edit History) on Qwen3-4B+SFT. A history length of 3 achieves the best balance, with 72.6% -do average accuracy and 89.4% -keep average accuracy. Short histories (e.g., 1) lack sufficient context for complex intents, while longer histories (e.g., 5+) introduce irrelevant noise, hindering accuracy.
RQ1 Answer about Edit History Length: Short histories may lack context, while long histories can introduce noise, both hindering accuracy. The results show that a history length of 3 is optimal.
@@
Seed-Coder-8B-Instrcut 71.30/14.9% 69.35/7.0% 69.91/8.8% 70.06/9.8% 70.16/10.1%
@@
Seed-Coder-8B-Instrcut+SFT 92.29/30.2% 92.85/34.0% 89.55/20.4% 89.54/22.0% 91.06/26.7%
@@
Seed-Coder-8B-Instrcut+SFT+DAPO 91.90/31.4% 93.21/34.2% 90.15/23.6% 90.08/26.0% 91.34/28.8%
@@
Seed-Coder-8B-Instrcut 2.6% 3.2% 2.0% 3.8% 2.9%
@@
Seed-Coder-8B-Instrcut+SFT 92.2% 86.0% 88.6% 89.8% 89.2%
@@
Seed-Coder-8B-Instrcut+SFT+DAPO 85.2% 80.0% 78.2% 83.4% 81.7%
@@ 4.2.2. RQ2 Code Edit Generation:
Impact of SFT: Table 4 and Table 5 present the performance of the NES-Edit Model across four programming languages: Java, Python, TypeScript, and TypeScript React. In the -do task, Zeta achieves the highest ES score of 85.94%, while Claude4 records the best EMR score of 23.1%. Following SFT fine-tuning, smaller models demonstrate improved ES and EMR scores. For instance, Qwen2.5-Coder-7B+SFT achieves an ES of 90.98% and EMR of 25.0%, Seed-Coder-8B-Instruct+SFT achieves an ES of 91.06% and EMR of 26.7%, and Qwen3-4B+SFT achieves an ES of 90.33% and EMR of 24.6%. In the -keep task, Zeta achieves the highest initial accuracy of 40.7%. After SFT fine-tuning, smaller models show significant improvements in accuracy, with Qwen2.5-Coder-7B+SFT reaching 91.0%, Seed-Coder-8B-Instruct+SFT achieving 89.2%, and Qwen3-4B+SFT attaining 88.5%.
@@
RQ2 Answer about SFT: The NES SFT dataset greatly improves model performance in code edit generation.
@@
Impact of DAPO: Smaller models trained on the DAPO dataset achieve moderate improvements in average code edit generation (-do) similarity scores. For example, Qwen2.5-Coder-7B improves from an ES of 90.98% to 91.13% and an EMR of 25.0% to 28.3%; Seed-Coder-8B-Instruct increases from an ES of 91.06% to 91.34% and an EMR of 26.7% to 28.8%; and Qwen3-4B improves from an ES of 90.33% to 91.36% and an EMR of 24.6% to 27.7%.
For the -keep task, Qwen2.5-Coder-7B experiences a decline in accuracy from 91.0% to 84.3%, Seed-Coder-8B-Instruct sees an decrease from 89.2% to 81.7%, and Qwen3-4B drops from 88.5% to 86.3%.
@@
The slight decrease in performance on the -keep task can be attributed to the same reason as with NES-Location: the DAPO-trained model tends to suggest location changes more frequently. Nevertheless, we deploy the DAPO-trained model in real-world scenarios due to its alignment with high-frequency practical use cases.
@@
RQ2 Answer about DAPO: The DAPO-trained model demonstrates improved similarity scores for the -do task, yet shows a minor drop in accuracy for the -keep task. However, we have chosen to adopt the model trained with DAPO, as it better aligns with the high-frequency practices observed in real-world software development.
@@
Impact of Unmodified Data: We investigated the impact of varying proportions of unmodified data in the training set, analyzing ratios from 10% to 60% (Table 6). The results show that higher proportions of unmodified data improve performance on the -keep test sets, while leading to a decline in performance on the -do test sets. After thorough analysis, a 20% proportion of unmodified data was determined to be optimal. As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
@@
RQ2 Answer about Unmodified Data: As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
Impact of Edit History Length: Table 7 illustrates the effect of varying the length of historical context (Max Edit History) on the performance of Qwen3-4B+SFT. Optimal results are achieved with a history length of 3, yielding a -do average similarity score of ES 90.33% and EMR 24.6%. Conversely, a history length of 7 results in the highest -keep average accuracy of 88.5%. This suggests that generating better code edits may require a more extensive historical trajectory.
@@
RQ2 Answer about Edit History Length: Generating better code edits may require a more extensive historical trajectory.
@@ 4.2.3. RQ3 Model Performance:
Table 8 presents the results of our inference speed analysis. The data reveals that both Prefix Caching (PC) and Speculative Decoding (SD) significantly enhance inference speed, with Speculative Decoding providing the largest improvement. This aligns with our initial hypothesis, confirming the suitability of Speculative Decoding for our specific use case.
When compared to Zeta, a quantized 7B model with the same inference optimizations, NES achieves superior input throughput, while maintaining equivalent output throughput. In our online environment, NES ultimately attained an average inference latency of 450ms, effectively reaching real-time processing capabilities and successfully meeting our performance objectives.
@@
Table 8. Comparison of input and output token processing speeds for different model configurations
RQ3 Answer: The Speculative Decoding (SD) optimization delivers the most significant improvement in inference input/output throughput, enabling NES to achieve an average inference latency of 450ms.
5. Design of Industrial Tools
In transitioning the NES framework from a research prototype to a scalable, industrial-grade solution, we faced the challenge of balancing model performance, inference latency, and deployment cost. To meet the stringent responsiveness demands of real-time code editing, we chose the Qwen3-4B model as the core engine for NES, post-train it with our SFT and DAPO datasets (see Section 3.3) for production deployment. This decision was based on several key considerations:
  • Trade-off between Model Size and Latency: Larger scale models, like 8B-parameter ones, have high computational needs and slower inference, making them unsuitable for real-time code editing’s low-latency requirements. The Qwen3-4B model, being smaller, offers faster inference, aligning better with our goal of achieving a "zero-latency" user experience in NES.
@@ 5. Design of Industrial Tools
  • Feasibility for Large-Scale Deployment: In industrial applications, ensuring service stability and scalability is critical. We deployed the Qwen3-4B+SFT+DAPO model across a configuration of NVIDIA L20 GPUs, leveraging the optimizations detailed in Section 3.5. This combination of model architecture and system-level enhancements reflects a design tailored to the practical constraints of large-scale deployment environments in real-world industrial scenarios. NES consistently achieved an average response time of under 450ms, delivering a seamless user experience while handling tens of thousands of daily calls. These results confirm the model’s effectiveness and reliability for industrial-grade applications.
@@ 6. Related Work
Transform-based Code Edits:
Several transformer-based approaches have been proposed to extend code generation solutions to code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Zhang et al., 2022a). CODIT (Chakraborty et al., 2018) utilizes AST-based operations for code edits but faces challenges with high computation and limited semantic understanding. SARGAM (Liu et al., 2023) introduces a three-stage framework (Search-Generate-Modify) combining retrieval, generation, and refinement for improved accuracy and efficiency. CodeT5 (Wang et al., 2021) is a versatile encoder-decoder model pre-trained for various code tasks. Although these methods differ in how they represent edits within deep learning models, they commonly frame the edit generation problem as a translation task, which often limits their ability to deliver an improved user experience.
General LLM-Powered Code Edits:
@@
In recent years, general LLMs have made significant progress in domains such as code generation (Chen et al., 2021; Zeng et al., 2025; Di et al., 2023; Jiang et al., 2024; Tian and Chen, 2023; AI, 2025), code completion (Liu et al., 2017; Alon et al., 2019; Feng et al., 2020; Liu et al., 2020; Chakraborty and Ray, 2021; Lu et al., 2022; Liang et al., 2024; Naznin and Mondal, 2023; Wang et al., 2024a; Wan et al., 2024; Wang et al., 2024b), and other code-related tasks (Liu et al., 2024b). Codex (Chen et al., 2021) has demonstrated the ability to generate entire functions and applications from complex natural language prompts. CodeFuse-13B (Di et al., 2023) is an open-sourced pre-trained Language Model (LLM) with 13 billion parameters designed for code-related tasks. The latest open-source model, Kimi K2 (AI, 2025), has achieved state-of-the-art (SOTA) results among open-source models on both the SWE-bench Verified (Aleithan et al., 2024) and ACeBench  (Chen et al., 2025), demonstrating its leading capabilities in code generation task. LLM-powered tools are well-suited for general-purpose coding tasks but often lack optimization and specialized techniques for code editing tasks, which demand stricter requirements such as low latency, high accuracy, and efficiency.
Specialized Code Edit:
@@
The advent of LLMs has spurred research in code editing, leading to the development of numerous approaches tailored specifically for code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Jin et al., 2025). InstructCoder (Li et al., 2023) introduces a dataset of 114,000 instruction-based triplets for LLM tuning, enhancing code editing accuracy for open-source models. However, it requires precise developer input, adding cognitive overhead. GRACE (Gupta et al., 2023) conditions LLMs on prior edits, significantly improving performance for edits distant from the current focus by capturing implicit developer intent. CoEdPilot (Liu et al., 2024a) combines historical edits, project context, and feedback to enhance code edit recommendations. It improves practicality but depends heavily on developer input and has high computational costs. Coeditor (Wei et al., 2024) uses repository-level changes and edit history to boost performance in single- and multi-turn editing. However, it may miss fine-grained edit patterns by focusing on repository diffs. Unlike these models that rely on developer prompts, NES takes a proactive approach by predicting edit locations and content. This enables a seamless, instruction-free workflow, fostering predictive collaboration.
7. Conclusion and Future Work
This paper presents NES, an LLM-powered framework for instruction-free and low-latency code editing assistance in IDEs. Unlike traditional approaches requiring explicit natural language instructions, NES leverages historical editing patterns to deliver efficient, context-aware support. It features two key components: the NES-Location Model, which predicts the next likely code modification location, and the NES-Edit Model, which generates contextually relevant code change suggestions based on editing history.
We evaluate NES using extensive real-world datasets from both open-source and industrial projects, highlighting its practical effectiveness. Experimental results show that the NES-Location Model achieves 75.6% accuracy in predicting the next edit location, while NES achieves 75.6% and 81.6% accuracy in two tasks of predicting next edit locations, alongside 91.36% ES and 27.7% EMR for intent-aligned edits, outperforming SOTA models. Also, our open-sourced SFT and DAPO datasets have been demonstrated to enhance the performance of open-source CodeLLMs.
Currently, NES operates within a single file, leveraging granular editing history in a localized context. Future work will enhance NES to handle multi-file code dependencies, enabling neural edit suggestions across interconnected files instead of single-file contexts.

### v3

@@ 4.2. Results
* Training large models on more data was infeasible because of the necessary low-latency performance and restrictive training resource constraints. Zeta was not evaluated because it lacks navigation functionality. CoEdPilot was evaluated without human instructions.
Table 3. Comparison of location prediction performance with varying max edit history lengths on the Qwen3-4B+SFT model
@@ 4.2.1. RQ1 NES-Location:
Before post-training, the best base model, Claude4, achieves 67.0% and 9.6% average accuracy on the -do and -keep tasks, respectively. Models with small size, such as Qwen2.5-Coder-7B, perform poorly, with only 10.1% and 3.7% on the -do and -keep tasks, respectively. NES SFT dataset significantly improves models with small size. For -do, Qwen2.5-Coder-7B jumps from 10.1% to 62.8%, Seed-Coder-8B-Instruct from 9.3% to 72.6%, and Qwen3-4B from 6.6% to 72.6%. For -keep, Qwen2.5-Coder-7B rises from 3.7% to 94.0%, Seed-Coder-8B-Instruct from 2.9% to 94.1%, and Qwen3-4B from 1.9% to 89.4%.
Limits. Due to constraints on low-latency requirements and training resources, large models were not trained on the SFT or DAPO dataset. Furthermore, since Zeta lacks navigation functionality, its results are excluded from Table 2. Finally, CoEdPilot’s reliance on explicit human inputs explains why its performance did not surpass that of other tools. A similar result was observed in the subsequent evaluation.
Impact of SFT: The NES SFT dataset significantly boosts model performance in predicting the next editing position, improving Qwen3-4B’s accuracy from 6.6% to 72.6% on the -do task and from 1.9% to 89.4% on the -keep task.
Training smaller models on the DAPO dataset results in a slight but notable improvement in average navigation prediction (-do) accuracy: Qwen2.5-Coder-7B improves from 62.8% to 65.6%, Seed-Coder-8B-Instruct increases from 72.6% to 77.7%, and Qwen3-4B rises from 72.6% to 75.6%. But, on the -keep task, Qwen2.5-Coder-7B declines from 94.0% to 86.5%, Seed-Coder-8B-Instruct declines from 94.1% to 85.0%, and Qwen3-4B drops from 89.4% to 81.6%.
The DAPO-trained model tends to suggest location changes more frequently. This is because changing is more frequent in the real-world development practices. We accept the minor drop in -keep performance in favor of improved -do accuracy. DAPO was chosen because predicting necessary jumps accurately matters more than avoiding suggestions in developers’ workflow.
Impact of DAPO: The DAPO-trained model tends to suggest location changes more frequently, leading to a slight drop in -keep task accuracy. However, the NES-Location model was chosen for its superior -do task performance, as location changes are more prevalent in practice.
Table 3 demonstrates the impact of historical context length (Max Edit History) on Qwen3-4B+SFT. A history length of 3 achieves the best balance, with 72.6% -do average accuracy and 89.4% -keep average accuracy. Short histories (e.g., 1) lack sufficient context for complex intents, while longer histories (e.g., 5+) introduce irrelevant noise, hindering accuracy.
Impact of Edit History Length: Short histories may lack context, while long histories can introduce noise, both hindering accuracy. The results show that a history length of 3 is optimal.
@@
Seed-Coder-8B-Instruct 71.30/14.9% 69.35/7.0% 69.91/8.8% 70.06/9.8% 70.16/10.1%
@@
Seed-Coder-8B-Instruct+SFT 92.29/30.2% 92.85/34.0% 89.55/20.4% 89.54/22.0% 91.06/26.7%
@@
Seed-Coder-8B-Instruct+SFT+DAPO 91.90/31.4% 93.21/34.2% 90.15/23.6% 90.08/26.0% 91.34/28.8%
@@
Seed-Coder-8B-Instruct 2.6% 3.2% 2.0% 3.8% 2.9%
@@
Seed-Coder-8B-Instruct+SFT 92.2% 86.0% 88.6% 89.8% 89.2%
@@
Seed-Coder-8B-Instruct+SFT+DAPO 85.2% 80.0% 78.2% 83.4% 81.7%
@@ 4.2.2. RQ2 Code Edit Generation:
Table 4 and Table 5 present the performance of the NES-Edit Model across four programming languages: Java, Python, TypeScript, and TypeScript React. In the -do task, Zeta achieves the highest ES score of 85.94%, while Claude4 records the best EMR score of 23.1%. Following SFT fine-tuning, smaller models demonstrate improved ES and EMR scores. For instance, Qwen2.5-Coder-7B+SFT achieves an ES of 90.98% and EMR of 25.0%, Seed-Coder-8B-Instruct+SFT achieves an ES of 91.06% and EMR of 26.7%, and Qwen3-4B+SFT achieves an ES of 90.33% and EMR of 24.6%. In the -keep task, Zeta achieves the highest initial accuracy of 40.7%. After SFT fine-tuning, smaller models show significant improvements in accuracy, with Qwen2.5-Coder-7B+SFT reaching 91.0%, Seed-Coder-8B-Instruct+SFT achieving 89.2%, and Qwen3-4B+SFT attaining 88.5%.
@@
Impact of SFT: The NES SFT dataset greatly improves model performance in code edit generation.
@@
Smaller models trained on the DAPO dataset achieve moderate improvements in average code edit generation (-do) similarity scores. For example, Qwen2.5-Coder-7B improves from an ES of 90.98% to 91.13% and an EMR of 25.0% to 28.3%; Seed-Coder-8B-Instruct increases from an ES of 91.06% to 91.34% and an EMR of 26.7% to 28.8%; and Qwen3-4B improves from an ES of 90.33% to 91.36% and an EMR of 24.6% to 27.7%.
@@
For the -keep task, Qwen2.5-Coder-7B experiences a decline in accuracy from 91.0% to 84.3%, Seed-Coder-8B-Instruct drops from 89.2% to 81.7%, and Qwen3-4B drops from 88.5% to 86.3%. The slight -keep drop is due to the DAPO-trained model suggesting changes more frequently, but we deploy it for its alignment with high-frequency practical use cases.
@@
Impact of DAPO: The DAPO-trained model demonstrates improved similarity scores for the -do task, yet shows a minor drop in accuracy for the -keep task. However, we have chosen to adopt the model trained with DAPO, as it better aligns with the high-frequency practices observed in real-world development.
@@
We investigated the impact of varying proportions of unmodified data in the training set, analyzing ratios from 10% to 60% (Table 6). The results show that higher proportions of unmodified data improve performance on the -keep test sets, while leading to a decline in performance on the -do test sets. After thorough analysis, a 20% proportion of unmodified data was determined to be optimal. As a result, we selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
@@
Impact of Unmodified Data: We selected the 20% unmodified data ratio as the standard configuration for all subsequent experiments.
Table 7 shows that a history length of 3 is optimal for the -do task (ES 90.33%, EMR 24.6%), while a length of 7 yields the best -keep accuracy (91.7%), suggesting better code edits may require longer historical trajectories.
@@
Impact of Edit History Length: Generating better code edits may require a more extensive historical trajectory.
@@ 4.2.3. RQ3 Model Performance:
Table 8 shows that Prefix Caching (PC) (Zheng et al., 2024), for reusing prompt computations, and Speculative Decoding (SD) (Chen et al., 2023; Leviathan et al., 2022), for faster token generation, significantly boost performance, with SD showing the largest gains. Evaluated on vLLM, NES demonstrates superior input throughput over Zeta, while a latency comparison with CoEdPilot was deemed infeasible, as its non-vLLM-based architecture would require a costly re-implementation for an equitable benchmark. These optimizations enable NES to achieve an average latency of 250ms in live deployment.
Table 8. Comparison of input and output token processing speeds for different model configurations.
@@
* CoEdPilot’s non-vLLM architecture is incompatible with PC and SD optimizations.
Answer to RQ3: The SD optimization delivers the most significant improvement in inference input/output throughput, enabling NES to achieve an average inference latency of 250ms.
5. Industrial Deployment
5.1. Design Rationale
In transitioning the NES framework to a scalable, industrial-grade solution, we faced the challenge of balancing model performance, inference latency, and deployment cost. To meet the stringent responsiveness demands of real-time code editing, we chose the Qwen3-4B model as the core engine, post-train it with our SFT and DAPO datasets (see Section 3.3) for production deployment. This decision was based on several key considerations:
  • Trade-off between Model Size and Latency: Larger scale models, like 8B-parameter ones, have high computational needs and slower inference, making them unsuitable for real-time code editing’s low-latency requirements. The Qwen3-4B model, being smaller, offers faster inference, aligning better with our goal of achieving a ”zero-latency” user experience in NES.
@@ 5.1. Design Rationale
  • Feasibility for Large-Scale Deployment: In industrial applications, ensuring service stability and scalability is critical. We deployed the Qwen3-4B+SFT+DAPO model across a configuration of NVIDIA L20 GPUs with the optimizations in Section 3.5, tailored for large-scale deployment. NES consistently achieved an average response time of under 250ms, delivering a seamless user experience while handling tens of thousands of daily calls. These results confirm the model’s effectiveness and reliability for industrial use.
5.2. Empirical Evaluation of User Experience
NES builds upon our prior work, PEACE (Ren et al., 2025), with repo-level program understanding (Zhang et al., 2025; Tao et al., 2025; Li et al., 2024), but its reliance on live user editing data—captured exclusively via our IDE plugin—mandates an in-the-wild evaluation with internal developers.
A direct comparison of acceptance rates is misleading due to fundamental design differences. Tools like CoEdPilot use a ”one-to-many” user-description-based model, which naturally produces high raw acceptance rates. In contrast, NES’s passive, ”one-by-one” suggestion model generates far more frequent interactions. To bridge this gap, we define an Equivalent Acceptance Rate (EAR) by multiplying the raw rate by the average number of lines accepted per event (1.84, based on 6 months of data).
This normalization shows NES’s effective rates for NES-Edit and NES-Location tasks are 43.44% and 51.55%, respectively. These results confirm that NES offers comparable quantitative performance while providing a superior, description-free user experience. This experiential advantage was further validated in a comparative user study against another tool, Zeta, where NES demonstrated similar positive outcomes to those detailed in Section 4.2. Notably, its effectiveness becomes more pronounced as system complexity increases; in scenarios involving large classes, intricate functions, or multi-application interactions within microservices, NES more reliably completes code editing tasks than alternative methods. We hypothesize that this superiority stems from its core design: by capturing the fine-grained user editing trajectory, NES implicitly learns the developer’s mental model and high-level understanding of the system’s architecture—a rich signal that is absent in methods relying solely on static code snapshots.
@@ 6. Related Work
Code Edits with LLMs
@@
Early transformer-based approaches extend code generation to editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Zhang et al., 2022b), including AST-based methods (Chakraborty et al., 2018), retrieval-augmented frameworks (Liu et al., 2023), and versatile encoder-decoder models such as CodeT5 (Wang et al., 2021). Recent LLMs have made significant progress in code generation (Chen et al., 2021; Zeng et al., 2025; Di et al., 2023; Codefuse et al., 2025; Jiang et al., 2024; Tian and Chen, 2023; AI, 2025), code completion (Liu et al., 2017; Alon et al., 2019; Feng et al., 2020; Liu et al., 2020; Chakraborty and Ray, 2021; Lu et al., 2022; Liang et al., 2024; Naznin and Mondal, 2023; Wang et al., 2024a; Wan et al., 2024; Wang et al., 2024b), and other code-related tasks (Liu et al., 2024b; Aleithan et al., 2024; Chen et al., 2025), with Codex (Chen et al., 2021) demonstrating the ability to generate entire functions and applications from complex natural language prompts. However, these tools often lack optimization for code editing’s strict requirements of low latency, high accuracy, and efficiency.
Specialized Code Edit
@@
The advent of LLMs has spurred research in code editing, leading to the development of numerous approaches tailored specifically for code editing tasks (Li et al., 2023; Gupta et al., 2023; Liu et al., 2024a; Wei et al., 2024; Jin et al., 2025; Wang et al., 2025). InstructCoder (Li et al., 2023) introduces a dataset of 114,000 instruction-based triplets for LLM tuning, enhancing code editing accuracy for open-source models. However, it requires precise developer input, adding cognitive overhead. CoEdPilot (Liu et al., 2024a) enhances code edit recommendations using historical edits and project context, but relies heavily on developer input and incurs high computational costs. Unlike prompt-based models, NES proactively predicts edit locations and content, enabling a seamless, instruction-free workflow.
7. Conclusion
We present NES, an LLM framework for proactive, instruction-free code editing that learns from user editing trajectories. It achieves state-of-the-art performance (91.36% Edit Similarity, 27.7% Exact Match), serves 20,000+ developers at Ant Group with under 250ms latency and Acceptance Rates of 43.44% and 51.55%, and publicly releases its training datasets.